please help me solve this word problem:
Jane leaves her house and drives in a straight line at a speed of 38 mph. Her speed on the return trip is 48 mph. The trip back home takes 15 minutes less than the outbound trip. What was Jane's total travel time in hours? Assume both trips are in a straight line. Give your answer to 2 decimal places.

Answer by Alan3354(69443)

Jane leaves her house and drives in a straight line at a speed of 38 mph. Her speed on the return trip is 48 mph. The trip back home takes 15 minutes less than the outbound trip. What was Jane's total travel time in hours? Assume both trips are in a straight line. Give your answer to 2 decimal places.
-------------------
15 mins = 1/4 hour
---
t = time at 38 mi/hr in hours
d = 38t
d = 48*(t- 1/4)
---
38t = 48t - 12
-10t = -12
t = 1.2 hours (at 38 mi/hr)
Total time = t + (t - 1/4)
= 2.15 hours